OpenStudy (anonymous):

solve the equation x^2-13x+30=0

OpenStudy (anonymous):

use quadratic formula

OpenStudy (abb0t):

are you trying to find the complimentary solution to a differential equation?

OpenStudy (austinl):

Or you could just factor. \(x^2-13x+30=0\)

OpenStudy (austinl):

That is an easy equation to factor out, it comes out beautifully.

OpenStudy (anonymous):

Yeah, just factor it out.

OpenStudy (anonymous):

When we have : x^2+bx+c=0 Use it : we have 2 answer : x1&x2 first find alfa : alfa : b^2-4ac x=-b+-\[\sqrt{alfa}\]

OpenStudy (abb0t):

You should be able to factor this out easily by now. Seeing as how you \(are\) in differential equations.

OpenStudy (austinl):

@abb0t : me? It is easy, the OP should be able to manage soundly by themselves. :P

OpenStudy (abb0t):

No. I was referring to @Roadhoue222

OpenStudy (anonymous):

the winner is Jaseem18

OpenStudy (austinl):

Ok, @austinL is confused right now...... whatever.

OpenStudy (anonymous):

lol